Suppose that Papa Bell, Inc.'s, equity is currently selling for $41 per share, with 3.6 million shares outstanding. The firm also has 8,000 bonds outstanding, which are selling at 95% of par. Assume Papa Bell was considering an active change to its capital structure so as to have a D/E of 0.5.

Which type of security (stocks or bonds) would the firm need to sell to accomplish this?

How much would it have to sell? (Enter your answer in dollars not in millions. Do not round intermediate calculations and round your final answer to 2 decimal places.)

Share Price: $41.00
Shares Outstanding: 3,600,000
Bonds Outstanding: 8,000
Bond Price (% of Par): 95%
Proposed New D/E Ratio: 0.50


Complete the following analysis. Do not hard code values in your calculations, and do not round intermediate calculations.

Current Equity Ratio:
Current Debt Ratio:
Current D/E Ratio:
Sell Bonds or Stock?:
New Debt Ratio:
Amount of Securities to Buy and Sell:
Answer #1
1Share   Outstanding 
36,00,000
2Share Price
$41.00
A.Current Value of   Equity(1*2)$14,76,00,000.00




3Bond Outstanding 
8000
4Bond Price (% of Par)
95%
5Par value of   Bond 
$1,000.00
6Bond Selling Price(4*5)$950.00
B.Current Value of   Bond (6*3)$76,00,000.00




7Total Assets (A+B)$15,52,00,000.00





Current Equity   RatioShareholder's   Equity/Total Assets = A/795.10%





Current Debt RatioValue of Debt/Total   Assets = B/74.9%

Current Debt/Equity   Ratio Debt/Equity = B/A0.0514905

1:Sell stocks and buy back bonds (This will increase the debt amount)

2:

Current value of debt$76,00,000
Value of equity$14,76,00,000
D/E Ratio0.0514905
Required D/E0.5
Debt required=Current debt/D/E Ratio$76,00,000/0.5 = $1,52,00,000
Hence Amount of Equity to be sold=Debt required-Current debt

$1,52,00,000 - $76,00,000 = $76,00,000

Value of Equity be = $14,76,00,000 - $76,00,000 = $14,00,00,000 

Value of Debt be = $1,52,00,000 

Total Assets = Value of Equity + Value of Debt = $14,00,00,000  + $1,52,00,000 = $15,52,00,000

New Debt Ratio = Value of Debt/Total Assets = $1,52,00,000/$15,52,00,000 = 0.097 or 9.79%